OpenStudy (anonymous):

sin (11pi/12) find exact value

hero (hero):

If you let pi = 180, you could put it degree form and figure it out from there.

OpenStudy (anonymous):

or you can think of it this way: \(\large sin(\frac{11\pi}{12})=sin(\frac{\frac{11\pi}{6}}{2}) \) .... Half-angle formula anyone?

hero (hero):

Now why would anyone want to do it that way?

hero (hero):

After simplifying I get \(\sin(165^{\circ})\)

OpenStudy (anonymous):

hmmm so it would be \[(\sqrt{6}+\sqrt{2})/4\]

hero (hero):

Which can be further broken down to \(\sin(120 + 45)\)

hero (hero):

Okay, so which hint did you use...mine or dpalnc?

OpenStudy (anonymous):

yours :O

hero (hero):

Awesome